Solve the problem as an equation.
You must show the addition/subtraction/division steps on both sides of the equation to receive full credit. You must show your work alegraically.
Kevin invested money in a savings account at a rate of 5% simple interest. After one year, he has $4032.00 in the account. How much did Kevin originally invest?

Answer by mananth(16946)

let the investment be x
5% *x = (4032-x)
multiply by 100
5x= 100(4032-x)
5x=403200-100x
5x+100x=403200-100x+100x
105x=403200
/105
x=403200/105
=3840